Suppose 100 fish are caught,tagged and released. Two weeks later 40 fish are caught, 5 of which are found have tags. Using this information,estimate the number of fish in the lake?
Answer by edjones(8007)

5/40=1/8 of the fish in the lake have been tagged.
100*8=800 fish in the lake.
.
Ed